OpenStudy (anonymous):

Please and thank youuuuu! What is the 23rd term of the arithmetic sequence where a1 = 7 and a5 = 21?

OpenStudy (cwrw238):

first find the common difference can u do that?

OpenStudy (anonymous):

Im not really sure what you mean?

OpenStudy (anonymous):

you have to find how far the first number is from the second one. that's the common difference.

OpenStudy (cwrw238):

arithmetc series can be written as a, a+d, a+ 2d, a + 3d, .... common difference d is added to each term to get the next term

OpenStudy (cwrw238):

a1 is first term and a5 is the fifth so 4 'd's have been added to get from a1 to a5 do you follow that ?

OpenStudy (anonymous):

got it

OpenStudy (cwrw238):

ok so if we work out a5 - a1 and divide by 4 we get d so 21-7 / 4 = 14/4 = 3.5 d = 3.5

OpenStudy (cwrw238):

general form for nth term = a1 + (n-1)d so 23rd term = 7 + (23-1)3.5 = 7 + 22*3.5

OpenStudy (anonymous):

84, thanks!

OpenStudy (cwrw238):

thats right - yw